A boat travel 30 miles upstream against the current in the same amount of time it take to travel 42 miles downstream with the current. If the rate of the current is 4MPH, what is the rate of the boat in still water?
Answer by checkley77(12844)

d=rt or t=d/r
t=30/(r-4)
t=42/(t+4)
Since the times are equal then:
30/(r-4)=42/(r+4)
42(r-4)=30(r+4)
42r-168=30r+120
42r-30r=120+168
12r=288
r=288/12
r=24 mph. is the boat's speed in still water.
Proof:
30/(24-4)=42/(24+4)
30/20=42/28
1.5=1.5